# LPO: Towards Accurate GUI Agent Interaction via Location Preference Optimization

### v2

@@ 5.1 Experimental Setup
Our agent is built upon the foundation model, Ovis2 8B [14]. During the SFT phase, we employ multiple inner datasets to equip the base model with GUI interaction capabilities. In the RL phase, we employ preference datasets from MMind2Web [3], AITZ [27], Omniact [9], OS-Genesis [20], Mug [10], and GUICourse [1] to optimize towards more accurate GUI interaction.
@@
To ensure a fair evaluation, we compare various preference optimization strategies using a single foundation model. Specifically, we select reward functions from UI-R1 [16] ($R_{UI\_R1}$), GUI-R1 [23] ($R_{GUI\_R1}$), and InfiGUI-R1 [13] ($R_{InfiGUI\_R1}$) as our baselines, each employing distinct preference optimization strategies, as illustrated in Figure 1.
@@
During the preference optimization, the training process lasted approximately 300 GPU hours, under the standard of the NVIDIA H100 GPU ***https://www.nvidia.com/en-sg/data-center/h100/.
@@
Following empirical insights from GRPO [19] and DAPO [25], we set the learning rate to $1\times 10^{-6}$ with a constant learning rate scheduler. Additionally, the lower clip range ($\epsilon_{1}$) is 0.2, while the upper clip range ($\epsilon_{2}$) is 0.28. The KL regularization hyperparameter ($\beta$) is adjusted to $1\times 10^{-4}$.
@@ 5.2 Offline Evaluation
We utilized the Multimodal Mind2Web [4] benchmark to assess the agent’s GUI interaction capabilities. This benchmark is specifically designed to create and evaluate agents’ capability to execute arbitrary tasks across various web environments.
@@
Table 1: Performance of GUI interaction on Multimodal Mind2Web [4]. We report Element Accuracy (Ele.Acc), Operation F1 (Op.F1) and Step Success Rate (Step SR). The best model is in-bold, and the second best is underlined.
Ele.Acc Op.F1 Step SR Ele.Acc Op.F1 Step SR Ele.Acc Op.F1 Step SR
After Supervised Fine-Tuning
Base Modal $60.3$ $57.4$ $38.2$ $60.7$ $56.9$ $38.4$ $63.8$ $58.5$ $40.7$
After Preference Optimization
+ $R_{UI\_R1}$ [16] $59.5$ $34.5$ $24.9$ $56.5$ $31.5$ $22.1$ $61.6$ $37.2$ $27.1$
+ $R_{GUI\_R1}$ [23] $62.5$ $\underline{71.6}$ $\underline{46.6}$ $61.2$ $67.6$ $43.5$ $65.0$ $71.1$ $47.9$
+ $R_{InfiGUI\_R1}$ [13] $\underline{62.6}$ $51.3$ $35.8$ $\underline{62.2}$ $49.5$ $34.4$ $\underline{65.1}$ $53.1$ $40.0$
GUI Grounding
To further evaluate the precise interaction capabilities of agents, we conduct evaluations to determine the effectiveness of preference optimization strategies on enhancing GUI grounding abilities. We employed VisualWebBench [12] and Screenspot V2 [22] as benchmarks, providing a broad spectrum of platforms to assess the capacity of GUI agents to accurately ground interaction locations.
VisualWebBench [12] offers a comprehensive evaluation framework by providing grounding-related tasks in website, element, and action. As shown in Table 2, our experimental results on this benchmark demonstrate that LPO consistently achieves SOTA performance and robustness across diverse environments. While GUI-R1 [23] shows enhanced WebQA performance, its effectiveness is restricted to particular scenarios and does not improve GUI grounding capabilities across multiple tasks substantially. In contrast, LPO shows clear superiority across various metrics, underscoring its robustness and SOTA performance in GUI grounding.
ScreenSpot V2 [22] provides a benchmark to directly locate text or icons/widgets across different device scenarios, including mobile, desktop, and web environments. As shown in Table 3, our experimental results indicate that LPO significantly and comprehensively enhances the visual localization capabilities of the base model across various terminal environments. While GUI-R1 [23] and InfiGUI-R1 [13] outperform LPO in a few specific tasks, their overall cross-scenario compatibility is considerably lower, resulting in overall performance that is only comparable to or slightly worse than the base model. In contrast, LPO improves upon the base model’s performance and achieves SOTA overall results compared to other baselines.
Table 2: Performance of GUI grounding on VisualWebBench [12]. ROUGE-L is used to measure the quality of the generated responses. WebQA is reported by style F1. For other multiple-choice tasks, we report accuracy. The best model is in-bold, and the second best is underlined.
@@
+ $R_{UI\_R1}$ [16] $23.2$ $78.0$ $69.2$ $96.5$ $96.8$ $96.0$ $91.2$ $78.7$
+ $R_{GUI\_R1}$ [23] $24.2$ $\mathbf{78.8}$ $69.3$ $96.6$ $96.8$ $96.4$ $91.6$ $78.8$
+ $R_{InfiGUI\_R1}$ [13] $23.7$ $75.9$ $69.2$ $96.3$ $96.8$ $96.7$ $\mathbf{91.7}$ $78.5$
@@
Table 3: Performance of GUI grounding on ScreenSpot V2 [22]. We report grounding accuracy in this table, determining correctness by whether a prediction falls within the ground truth bounding box. The best model is in-bold, and the second best is underlined.
@@
+ $R_{UI\_R1}$ [16] $97.5$ $77.7$ $93.8$ $82.1$ $94.0$ $84.2$ $88.2$
+ $R_{GUI\_R1}$ [23] $97.5$ $77.7$ $94.8$ $84.2$ $93.5$ $\mathbf{84.7}$ $88.7$
+ $R_{InfiGUI\_R1}$ [13] $\mathbf{98.2}$ $80.0$ $95.3$ $86.0$ $93.5$ $83.2$ $89.5$
@@
5.3 Online Evaluation
To thoroughly assess the applicability of our preference optimization strategy in real-world scenarios, we conducted online evaluations to directly measure the performance of the GUI Agent in dynamic online environments.
We utilized WebVoyager [7] as our benchmark, performing online evaluations on nine accessible websites: Amazon, Apple, Arxiv, BBC News, Coursera, GitHub, Hugging Face, Wolfram Alpha, and ESPN. Other websites were unavailable due to network issues (Google Search and Google Map), timeliness (Booking, Google Flights), and anti-scraping measures (Allrecipes, Cambridge Dictionary).
Table 4: Performance of online evaluation on WebVoyager [7]. We report the Task Success Rate in the table. The best model is in-bold, and the second best is underlined.
Amazon Apple ArXiv BBC News Coursera
@@ 5.3 Online Evaluation
Base Model $\underline{40.0}$ $58.1$ $53.4$ $38.0$ $54.7$
@@
+ $R_{UI\_R1}$ [16] $12.2$ $41.8$ $51.1$ $30.9$ $45.2$
+ $R_{GUI\_R1}$ [23] $35.0$ $37.2$ $27.9$ $33.3$ $57.1$
+ $R_{InfiGUI\_R1}$ [13] $\mathbf{51.2}$ $51.1$ $55.8$ $\mathbf{59.5}$ $69.0$
+ LPO (Ours) $\mathbf{51.2}$ $\mathbf{60.5}$ $\mathbf{64.3}$ $54.7$ $\mathbf{71.4}$
Github Huggingface Wolfram Alpha ESPN Overall
After Supervised Fine-Tuning
Base Model $\mathbf{65.8}$ $33.3$ $56.5$ $\mathbf{41.8}$ $48.0$
After Preference Optimization
+ $R_{UI\_R1}$ [16] $58.3$ $\mathbf{51.1}$ $63.0$ $27.9$ $47.3$
+ $R_{GUI\_R1}$ [23] $50.0$ $35.0$ $56.5$ $15.9$ $37.5$
+ $R_{InfiGUI\_R1}$ [13] $53.6$ $43.6$ $\mathbf{65.9}$ $\mathbf{41.8}$ $54.1$
+ LPO (Ours) $56.1$ $47.5$ $57.5$ $38.6$ $\mathbf{57.6}$
5.4 Ablation Study
We conduct ablation experiments on our two rewards proposed in this paper and analyze their impact on the overall performance in Table 5 w/o $r_{d}$ and w/o $r_{w}$.
Table 5: Performance of ablation study on Multimodal Mind2Web [4]. The best model is in-bold, and the second best is underlined.
@@ 5.4 Ablation Study
w/o $r_{d}$ $56.7$ $\underline{74.6}$ $42.3$ $56.3$ $69.7$ $40.9$ $61.2$ $\underline{73.1}$ $45.6$
w/o $r_{w}$ $\underline{62.7}$ $71.7$ $\underline{46.4}$ $\underline{61.6}$ $\underline{70.3}$ $\underline{44.1}$ $\underline{64.2}$ $71.9$ $\underline{47.6}$
@@
To demonstrate the efficacy of the window-based information density reward $r_{w}$, we compare the performance of our optimization strategy with and without $r_{w}$. As shown in Table 5 (w/o $r_{w}$), the absence of $r_{w}$ leads to a decline in operational accuracy, underscoring the importance of focusing on high-density informational areas to enhance the agent’s decisiveness and effectiveness in GUI agent interaction.
@@
To validate the effectiveness of the dynamic location reward $r_{d}$, we similarly compared our performance with and without $r_{d}$. As indicated in Table 5 (w/o $r_{d}$), the exclusion of $r_{d}$ result in a significant reduction in element accuracy due to the absence of spatial relationship. This highlights the substantial impact of dynamic location reward on GUI spatial optimization. Additionally, the success rate per action and operational correctness also declined, demonstrating the critical role of location information in action decision-making.
6 Limitations
Dependence on Extensive High-Precision Location Datasets
@@ 6 Limitations
Significant Computational Overhead
Training the LPO approach demands considerable computational power due to its complex integration of locational data and dynamic reward mechanisms. This high computational requirement can hinder real-time application scenarios and limit accessibility to users with less advanced computing resources.
7 Conclusion
In this paper, we delved into the challenge of achieving high-accuracy interactions for autonomous agents in GUI. We highlighted the inherent difficulties associated with spatial localization when conventional supervised fine-tuning methods are employed, which often fall short in providing the required precision for such tasks. Recognizing these limitations, we proposed a novel solution: Location Preference Optimization (LPO). This approach is designed to refine interaction accuracy by utilizing locational data to inform and optimize interaction preferences, thus addressing the shortcomings of existing methodologies.
LPO significantly improves GUI agents’ interaction capabilities, demonstrating superior performance in both offline benchmarks and online evaluations. This advancement sets a new standard for precision in GUI interactions and lays the groundwork for more intelligent and adaptive systems, offering a promising direction for future developments in complex interface interactions.
Appendix A Social Impact
@@
Appendix B Future Work

### v3

@@ 5.1 Experimental Setup
Our agent is built upon the foundation model, Ovis2 8B Lu et al. (2024b). During the SFT phase, we employ multiple inner datasets to equip the base model with GUI interaction capabilities. In the RL phase, we employ preference datasets from MMind2Web Deng et al. (2023a), AITZ Zhang et al. (2024b), Omniact Kapoor et al. (2024), OS-Genesis Sun et al. (2024), Mug Li et al. (2022), and GUICourse Chen et al. (2024) to optimize towards more accurate GUI interaction.
@@
To ensure a fair evaluation, we compare various preference optimization strategies using a single foundation model. Specifically, we select reward functions from UI-R1 Lu et al. (2025) ($R_{\text{UI-R1}}$), GUI-R1 Xia and Luo (2025) ($R_{\text{GUI-R1}}$), and InfiGUI-R1 Liu et al. (2025) ($R_{\text{InfiGUI-R1}}$) as our baselines, each employing distinct preference optimization strategies, as illustrated in Figure 1.
@@
During the preference optimization, the training process lasted approximately 300 GPU hours, under the standard of the NVIDIA H100 GPU.
@@
Following empirical insights from GRPO Shao et al. (2024) and DAPO Yu et al. (2025), we set the learning rate to $1\times 10^{-6}$ with a constant learning rate scheduler. Additionally, the lower clip range ($\epsilon_{1}$) is 0.2, while the upper clip range ($\epsilon_{2}$) is 0.28. The KL regularization hyperparameter ($\beta$) is adjusted to $1\times 10^{-4}$.
@@ 5.2 Offline Evaluation
We utilized the Multimodal Mind2Web Deng et al. (2023b) benchmark to assess the agent’s GUI interaction capabilities. This benchmark is specifically designed to create and evaluate agents’ capability to execute arbitrary tasks across various web environments.
@@
Table 2: Performance of GUI grounding on VisualWebBench Liu et al. (2024). ROUGE-L is used to measure the quality of the generated responses. WebQA is reported by style F1. For other multiple-choice tasks, we report accuracy. The best model is in-bold, and the second best is underlined.
@@
+ $R_{\text{UI-R1}}$ Lu et al. (2025) $23.2$ $78.0$ $69.2$ $96.5$ $96.8$ $96.0$ $91.2$ $78.7$
+ $R_{\text{GUI-R1}}$ Xia and Luo (2025) $24.2$ $\mathbf{78.8}$ $69.3$ $96.6$ $96.8$ $96.4$ $91.6$ $78.8$
+ $R_{\text{InfiGUI-R1}}$ Liu et al. (2025) $23.7$ $75.9$ $69.2$ $96.3$ $96.8$ $96.7$ $\mathbf{91.7}$ $78.5$
@@
Table 3: Performance of GUI grounding on ScreenSpot V2 Wu et al. (2024). We report grounding accuracy in this table, determining correctness by whether a prediction falls within the ground truth bounding box. The best model is in-bold, and the second best is underlined.
@@
+ $R_{\text{UI-R1}}$ Lu et al. (2025) $97.5$ $77.7$ $93.8$ $82.1$ $94.0$ $84.2$ $88.2$
+ $R_{\text{GUI-R1}}$ Xia and Luo (2025) $97.5$ $77.7$ $94.8$ $84.2$ $93.5$ $\mathbf{84.7}$ $88.7$
+ $R_{\text{InfiGUI-R1}}$ Liu et al. (2025) $\mathbf{98.2}$ $80.0$ $95.3$ $86.0$ $93.5$ $83.2$ $89.5$
@@
GUI Grounding
To further evaluate the precise interaction capabilities of agents, we conduct evaluations to determine the effectiveness of preference optimization strategies on enhancing GUI grounding abilities. We employed VisualWebBench Liu et al. (2024) and Screenspot V2 Wu et al. (2024) as benchmarks, providing a broad spectrum of platforms to assess the capacity of GUI agents to accurately ground interaction locations.
VisualWebBench Liu et al. (2024) offers a comprehensive evaluation framework by providing grounding-related tasks in website, element, and action. As shown in Table 2, our experimental results on this benchmark demonstrate that LPO consistently achieves SOTA performance and robustness across diverse environments. While GUI-R1 Xia and Luo (2025) shows enhanced WebQA performance, its effectiveness is restricted to particular scenarios and does not improve GUI grounding capabilities across multiple tasks substantially. In contrast, LPO shows clear superiority across various metrics, underscoring its robustness and SOTA performance in GUI grounding.
ScreenSpot V2 Wu et al. (2024) provides a benchmark to directly locate text or icons/widgets across different device scenarios, including mobile, desktop, and web environments. As shown in Table 3, our experimental results indicate that LPO significantly and comprehensively enhances the visual localization capabilities of the base model across various terminal environments. While GUI-R1 Xia and Luo (2025) and InfiGUI-R1 Liu et al. (2025) outperform LPO in a few specific tasks, their overall cross-scenario compatibility is considerably lower, resulting in overall performance that is only comparable to or slightly worse than the base model. In contrast, LPO improves upon the base model’s performance and achieves SOTA overall results compared to other baselines.
Table 4: Performance of online evaluation on WebVoyager He et al. (2024). We report the Task Success Rate in the table. The best model is in-bold, and the second best is underlined.
Method Amazon Apple ArXiv BBC News Coursera Github Huggingface Wolfram Alpha ESPN Overall
@@
Base Model $\underline{40.0}$ $\underline{58.1}$ $53.4$ $38.0$ $54.7$ $\mathbf{65.8}$ $33.3$ $56.5$ $\mathbf{41.8}$ $48.0$
@@
+ $R_{\text{UI-R1}}$ Lu et al. (2025) $12.2$ $41.8$ $51.1$ $30.9$ $45.2$ $\underline{58.3}$ $\mathbf{51.1}$ $63.0$ $27.9$ $47.3$
+ $R_{\text{GUI-R1}}$ Xia and Luo (2025) $35.0$ $37.2$ $27.9$ $33.3$ $\underline{57.1}$ $50.0$ $35.0$ $56.5$ $15.9$ $37.5$
+ $R_{\text{InfiGUI-R1}}$ Liu et al. (2025) $\mathbf{51.2}$ $51.1$ $\underline{55.8}$ $\mathbf{59.5}$ $\underline{69.0}$ $53.6$ $43.6$ $\mathbf{65.9}$ $\mathbf{41.8}$ $\underline{54.1}$
+ LPO (Ours) $\mathbf{51.2}$ $\mathbf{60.5}$ $\mathbf{64.3}$ $\underline{54.7}$ $\mathbf{71.4}$ $56.1$ $\underline{47.5}$ $\underline{57.5}$ $\underline{38.6}$ $\mathbf{57.6}$
Table 5: Performance of ablation study on Multimodal Mind2Web Deng et al. (2023b). The best model is in-bold, and the second best is underlined.
@@
w/o $r_{\mathrm{d}}$ $56.7$ $\underline{74.6}$ $42.3$ $56.3$ $69.7$ $40.9$ $61.2$ $\underline{73.1}$ $45.6$
w/o $r_{\mathrm{w}}$ $\underline{62.7}$ $71.7$ $\underline{46.4}$ $\underline{61.6}$ $\underline{70.3}$ $\underline{44.1}$ $\underline{64.2}$ $71.9$ $\underline{47.6}$
@@
5.3 Online Evaluation
To thoroughly assess the applicability of our preference optimization strategy in real-world scenarios, we conducted online evaluations to directly measure the performance of the GUI Agent in dynamic online environments. We utilized WebVoyager He et al. (2024) as our benchmark, performing online evaluations on nine accessible websites: Amazon, Apple, Arxiv, BBC News, Coursera, GitHub, Hugging Face, Wolfram Alpha, and ESPN. Other websites were unavailable due to network issues (Google Search and Google Map), timeliness (Booking, Google Flights), and anti-scraping measures (Allrecipes, Cambridge Dictionary).
5.4 Ablation Study
We conduct ablation experiments on our two rewards proposed in this paper and analyze their impact on the overall performance in Table 5 w/o $r_{\mathrm{d}}$and w/o $r_{\mathrm{w}}$.
@@ 5.4 Ablation Study
To demonstrate the efficacy of the window-based information density reward $r_{\mathrm{w}}$, we compare the performance of our optimization strategy with and without $r_{\mathrm{w}}$. As shown in Table 5 (w/o $r_{\mathrm{w}}$), the absence of $r_{\mathrm{w}}$ leads to a decline in operational accuracy, underscoring the importance of focusing on high-density informational areas to enhance the agent’s decisiveness and effectiveness in GUI agent interaction.
@@
To validate the effectiveness of the dynamic location reward $r_{\mathrm{d}}$, we similarly compared our performance with and without $r_{\mathrm{d}}$. As indicated in Table 5 (w/o $r_{\mathrm{d}}$), the exclusion of $r_{\mathrm{d}}$ results in a significant reduction in element accuracy due to the absence of spatial relationships. This highlights the substantial impact of dynamic location reward on GUI spatial optimization. Additionally, the success rate per action and operational correctness also declined, demonstrating the critical role of location information in action decision-making.
6 Conclusion
In this paper, we delve into the challenge of achieving high-accuracy interactions for autonomous agents in GUI. We propose a novel solution: Location Preference Optimization (LPO). This approach is designed to refine interaction accuracy by utilizing locational data to inform and optimize interaction preferences. LPO significantly improves GUI agents’ interaction capabilities, demonstrating superior performance in both offline benchmarks and online evaluations. This advancement lays the groundwork for more intelligent and adaptive systems, offering a promising direction for future developments in complex interface interactions.
Limitations
@@ 6 Conclusion
Besides, training the LPO approach demands considerable computational power due to its complex integration of locational data and dynamic reward mechanisms. This high computational requirement can hinder real-time application scenarios and limit accessibility to users with less advanced computing resources.
Acknowledgments
The work was supported by the Research Grants Council of HKSAR under grant number AoE/E-601/24-N.
Appendix
This appendix provides supplementary details on training cost, image resizing and distance normalization, additional visual grounding comparisons, and hyperparameter sensitivity analysis, supporting the main experimental findings of LPO.
Appendix A Training Cost
The main text reports $\sim$300 GPU hours on a single NVIDIA H100; with 8-way parallelization, wall-clock training is typically on the order of $\sim$40 hours for our 8B setup—commensurate with large-scale preference optimization. Table 6 compares approximate single-GPU hours against Multimodal Mind2Web Cross-Task Step SR when methods share the same backbone and preference-data mixture.
Table 6: Training cost vs. Mind2Web Cross-Task Step SR (approximate single-H100 hours).
[…]
+ $R_{\text{UI-R1}}$ $\sim$180
+ $R_{\text{GUI-R1}}$ $\sim$220
+ $R_{\text{InfiGUI-R1}}$ $\sim$260
+ LPO (Ours) $\sim$300
Appendix B Image Resizing and $d_{\max}$
All RGB observations are resized so that the longest edge is 1000 pixels (aspect ratio preserved). Thus coordinate distances live in a bounded box whose diameter is at most $1000\sqrt{2}$ pixels; setting $d_{\max}=1000$ in $r_{\mathrm{d}}$ defines a stable normalization in this space. Deployments that preserve native resolution should consider adaptive normalization (e.g., by image diagonal or reference bounding-box size); we discuss this under Future Work below.
Appendix C Visual Grounding Comparison
Table 7 compares ScreenSpot V2 accuracy against representative grounding models (same point-in-box criterion as the main paper). Training data and objectives differ across rows; the comparison positions LPO relative to dedicated grounding systems.
Table 7: Representative visual grounding methods on ScreenSpot V2 (%). “Avg.” matches the main paper’s aggregation.
Model Mob.-Txt Mob.-Icon Desk.-Txt Desk.-Icon Web-Txt Web-Icon Avg.
Proprietary
Operator 47.3 41.5 90.2 80.3 92.8 84.3 70.5
GPT-4o + OmniParser-v2 95.5 74.6 92.3 60.9 88.0 59.6 80.7
General open-source
Qwen2.5-VL-3B 93.4 73.5 88.1 58.6 88.0 71.4 80.9
Qwen2.5-VL-7B 97.6 87.2 90.2 74.2 93.2 81.3 88.8
GUI-specific (SFT)
SeeClick-9.6B 78.4 50.7 70.1 29.3 55.2 32.5 55.1
Magma-8B 62.8 53.4 80.0 57.9 67.5 47.3 61.5
OS-Atlas-4B 87.2 59.7 72.7 46.4 85.9 63.1 71.9
UI-TARS-2B 95.2 79.1 90.7 68.6 87.2 78.3 84.7
OS-Atlas-7B 95.2 75.8 90.7 63.6 90.6 77.3 84.1
Aguvis-7B 95.5 77.3 95.4 77.9 91.0 72.4 86.0
UGround-V1-7B 95.0 83.3 95.0 77.8 92.1 77.2 87.6
UI-TARS-72B 94.8 86.3 91.2 87.9 91.5 87.7 90.3
GUI-specific (RL)
SE-GUI-7B 90.3
LPO-8B (Ours) 97.9 82.9 95.9 86.4 95.6 84.2 90.5
Appendix D Hyperparameter Sensitivity
Table 8 varies group size $G$, clip ranges $(\epsilon_{1},\epsilon_{2})$, and KL coefficient $\beta$ on Mind2Web Cross-Task Step SR. Our default $(G,\epsilon_{1},\epsilon_{2},\beta)=(16,0.2,0.28,10^{-4})$ achieves the best trade-off in this sweep.
Table 8: Hyperparameter sensitivity (Mind2Web Cross-Task Step SR, %).
$G$ $\epsilon_{1}$ $\epsilon_{2}$ $\beta$ Step SR
8 0.2 0.28 $10^{-4}$ 48.1
16 0.2 0.28 $10^{-4}$ 49.5
16 0.1 0.2 $10^{-4}$ 48.7
16 0.3 0.4 $10^{-4}$ 48.9
16 0.2 0.28 $5{\times}10^{-5}$ 49.2
16 0.2 0.28 $5{\times}10^{-4}$ 47.8
Appendix E Social Impact
@@
Appendix F Future Work
@@
Adaptive Distance Normalization and Element Geometry
Beyond fixed $d_{\max}$, distance rewards could incorporate image diagonal, viewport scale, or per-element bounding-box extent so that pixel errors are measured relative to control size to reduce mismatch between small icons and large containers.
Semantic and Structure-Aware Entropy
Pixel entropy may over-weight visually busy but non-interactive regions. Combining low-level entropy with semantic segmentation, DOM structure, or saliency from UI models could improve alignment with true click likelihood.
Sample-Efficient and Stable RL
Exploring variance-reduction, shorter rollouts, or offline preference learning may lower the $\sim$300 GPU-hour budget while preserving the benefits of dense spatial rewards.